sorry if i keep sending so many problems but i need help on some problems i dont get multi step equations so yeah here are the problems,8x-5=4x+4-2x;15-3y=y+13+y;3w-1-4w=4-2w
Answer by stanbon(48568)

8x-5=4x+4-2x
8x-5 = 2x+4
6x = 9
x = 3/2
--------------
15-3y=y+13+y
15-3y = 2y+13
Add 3y to both sides to get:
15 = 5y+13
Subtract 13 from both sides to get:
5y= 2
Divide both sides by 5 to get:
y = 2/5
---------------
3w-1-4w=4-2w
-w-1 = 4-2w
Add 2w to both sides to get:
w-1 = 4
Add 1 to both sides to get:
w = 5
==================
Cheers,
Stan H.